OpenStudy (kitkat16):

Can someone correct this for me? PLEASE the weight of 1000 men normal distribution with a mean of 150 pounds and a standard deviation of 15. What the number of men weighing more than 165? Whats the number of men weighing less than 135? Is it 37% 226 men weigh more than 165 and 85 men weigh less than 135? The more I do this the more confused I get. I thought the percent had to fall between 68, 95, 99.7. Please HELP!

OpenStudy (amistre64):

165 = 150 + 15 = 1sd above the mean. X > 165 is the probability (the area under the curve) to the right of 1sd from the mean

OpenStudy (amistre64):

the empirical rule, for approximations, gives us that 68% of the data is centered about the mean within +-1 sd of the mean ... 1sd to the left of it, and 1sd to the right of it. 34% to the left of the mean and 34% to the right of the mean the mean itself sits at 50%, so the total amount of data to the left of the mean is 50%, and the total amount of data to the right of the mean is 50% what does this tell us? well, out of 50%, 34% is used up: 50 - 34 = 16 there is 16% of that data above 1sd to the right. therefore, 16% of the 1000 men, are expected to have a weight of more than 165

OpenStudy (kitkat16):

So 160 men weighed more than 165

OpenStudy (amistre64):

since 135 is 15 from 150 .... we have the same construction but in the opposite direction 135 is -1sd from the mean ... and is expected to have the same probability as 1sd above the mean so the solutions are the same.

OpenStudy (amistre64):

yes ... 16*10 is 160 is an approximation if we are to use the emipirical rule.

OpenStudy (amistre64):

http://www.tc3.edu/instruct/sbrown/swt/pic/chap03_empirical.jpg this gives the overall view of what we are looking at.

OpenStudy (kitkat16):

so is it 113 weigh less than 135?

OpenStudy (amistre64):

of course not 150 - 135 = 15, and 1sd = 15 since the normal distribution is exactly alike on the left as on the right ... the process is going to work out exactly the same ... and we are going to have about 16% tht weigh less than 135

OpenStudy (amistre64):

165 - 150 = 15; 1sd 150 - 135 = 15; 1sd they are the same ...

OpenStudy (kitkat16):

wait so its that easy?

OpenStudy (kitkat16):

wouldn't it be 165+15=180 and 135-15=120